DISPERSED FLUORESCENCE SPECTROSCOPY OF H_{2}CO IN A FREE-JET EXPANSION

\underline{KRISTIN K. ELLSWORTH} and WILLIAM F. POLIK, Department of Chemistry, Hope College, Holland, MI 49423.

High resolution dispersed fluorescence (DF) spectra of the excited vibrational levels in formaldehyde have been recorded to describe its S_{0} potential energy surface. Formaldehyde was cooled in a free-jet expansion and excited with a laser to the 4^{1}, 4^{0}, 3^{1}4^{1} and 5^{1} levels in S_{1}. The resulting fluorescence was dispersed with a monochromator to yield S_{0} vibrational spectra. Over 268 vibrational transitions have been assigned and fit to a polyad model. The model includes harmonic and anharmonic terms, and also accounts for resonances among vibrational levels. The results are compared to ab initio calculations.